OpenStudy (anonymous):

Can you solve log(4k-5)=log(2k-1)? Also, can you show the steps on how to solve an equation like this? Thanks.

OpenStudy (mathstudent55):

If \( \log a= \log b\) then \(a = b\).

OpenStudy (anonymous):

(4k-5) and (2k-1).

OpenStudy (mathstudent55):

Make them equal and solve for k.

OpenStudy (anonymous):

So, it'd be 4k-2k=-1+5?

OpenStudy (mathstudent55):

Correct. Now combine like terms on both sides.

OpenStudy (anonymous):

Is it 3 then? My friend got 2 so i was confused.

OpenStudy (anonymous):

Thanks anyways.

OpenStudy (mathstudent55):

Sorry, I lost the connection for a while. I'm back now.

OpenStudy (mathstudent55):

This line is correct. 4k-2k=-1+5 Now subtract 4k - 2k on the left side and add -1 + 5 on the right side: 2k = 4 Finally divide both sides by 2 k = 2

OpenStudy (mathstudent55):

You're friend is correct. The answer is k = 2.